# The Rate-Distortion-Perception-Classification Tradeoff: Joint Source Coding and Modulation via Inverse-Domain GANs

### v1

@@ Appendix B Proof of Lemma 3
=:$=\mathrel{\mathop{\ordinarycolon}}W_{1}\Big{(}p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{0}\Big{)}\cdot p_{0}+W_{1}\Big{(}p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{1}\Big{)}\cdot p_{1}$ (73)
@@
In (72), we first conditioned on $H_{0}$ and $H_{1}$, and then used the subadditivity of the supremum. The inequality is due to using different variables $f$ and $\widehat{f}$. From (72) to (73), we defined the Wasserstein-$1$ conditional on an event. From (73) to (74), we used the fact that $W_{p}(\cdot,\,\cdot)\leq W_{q}(\cdot,\,\cdot)$ whenever $p\leq q$; see [39, Remark 6.6]. From (74) to (75), we applied (28) to the models in (1) and (IV-B). And in the last step, we used the triangular inequality. ∎

### v2

@@ Appendix B Proof of Lemma 3
For convenience, we restate Lemma 3:
Lemma.
Let $p_{\bm{X}}$ (resp. $p_{\widehat{\bm{X}}}$) be a Gaussian mixture model following (12) [resp. (13)], in which the probability of hypothesis $H_{0}$ is $p_{0}$ and of hypothesis $H_{1}$ is $p_{1}=1-p_{0}$. Then,
$W_{1}(p_{\bm{X}},\,p_{\widehat{\bm{X}}})\leq\big{\|}\widehat{\bm{\Sigma}}^{\frac{1}{2}}-\bm{I_{n}}\big{\|}_{F}+\big{\|}\bm{D}\bm{E}\bm{c_{n}}-\bm{c_{n}}\big{\|}_{2}\cdot p_{1}\,.$
@@
$=:W_{1}\Big{(}p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{0}\Big{)}\cdot p_{0}+W_{1}\Big{(}p_{\bm{X}},\,p_{\widehat{\bm{X}}}\,|\,H_{1}\Big{)}\cdot p_{1}$ (76)
@@
In (75), we first conditioned on $H_{0}$ and $H_{1}$, and then used the subadditivity of the supremum. The inequality is due to using different variables $f$ and $\widehat{f}$. From (75) to (76), we defined the Wasserstein-$1$ conditional on an event. From (76) to (77), we used the fact that $W_{p}(\cdot,\,\cdot)\leq W_{q}(\cdot,\,\cdot)$ whenever $p\leq q$; see [Villani09-OptimalTransport, Remark 6.6]. From (77) to (78), we applied (23) to the models in (12) and (13). And in the last step, we used the triangular inequality. ∎